Grade 11Most Scoring Topics in IIT JEE

Jones covered a distance of 50 miles on his first trip. On a later trip he traveled 300 miles while going three times as fast. His new time compared with the old time was:

(a) three times as much, (b) twice as much, (c) the same, (d) half as much, (e) a third as much

User
9 Years agoGrade 11

2 Answers

User
9 Years ago
Let speed of the 1st trip x miles / hr. and speed of the 2nd trip 3x / hr.

We know that

Speed = Distance/Time.

Or, Time = Distance/Speed.

So, times taken to covered a distance of 50 miles on his first trip = 50/x hr.

And times taken to covered a distance of 300 miles on his later trip = 300/3x hr.

= 100/x hr.

So we can clearly see that his new time compared with the old time was: twice as much.

Answer: (b)
MOHIT GARG
9 Years ago
Let speed of the 1st trip Y miles / hr. and speed of the 2nd trip 3Y / hr.

We know that
SPEED-DISTANCE/TIME

Time = Distance/Speed.

 Time taken to cover a distance of 50 miles = 50/Y hr.

 Time taken to cover a distance of 300 miles = 300/3Y hr.= 100/Y hr.

COMPARING WE GET=100/Y*(Y/50)=2
 
OR NEW TIME IS TWICE THE FIRST ONE ..

Answer: (b)